### v1

# BARI-TH/25-772

Abstract
We consider an abelian extension of the Standard Model (SM) comprising a new gauge group $U(1)^{\prime}$, with the neutral gauge boson $Z^{\prime}$ having flavour violating couplings to quarks and leptons. The fermion content is the same as in SM except for the addition of three right-handed neutrinos. The model, proposed in [1], describes the couplings of $Z^{\prime}$ to fermions in terms of three rational parameters $\epsilon_{1,2,3}$ that sum to zero imposing the cancellation of the gauge anomalies. Each $\epsilon_{i}$ is common to all fermions in a generation, a feature producing correlations among quark and lepton observables. We focus on $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transitions for the lepton flavour conserving $\ell_{1}=\ell_{2}$ and lepton flavour violating case $\ell_{1}\neq\ell_{2}$. Small deviations with respect to the SM predictions are found in the first case, which reflects a feature of the model where quark and lepton sectors prevent each other to manifest large discrepancies with respect to SM. We investigate the correlations with the flavour violating leptonic decays $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$. The experimental upper bound on the branching ratio of $\mu^{-}\to e^{-}\gamma$ constrains the range for the lepton flavour violating $B_{d,s}$ decays, which however are predicted to be within the reach of the planned new facilities.
1 Introduction
The Standard Model of particle physics (SM) has been directly probed up to the energies accessible by the Large Hadron Collider and, within the present theoretical and experimental accuracy, the tests have been successful: The model provides us with a careful description of the interactions among the basic constituents of matter, excluding gravity. However, there are well known and broadly shared arguments to assert that this is not the ultimate theory for the fundamental interactions. Apart from the missing description of gravity, this model does not give any insight, for example, on the large number of independent parameters and the hierarchical structure of some of them, the regularities in the fermion sector, e.g. the organisation in generations, the texture of the fermion mixing matrices, and so on. A theory extended at higher energies, able to face at least some of the unsolved issues is actively sought for. The Standard Model would be its low-energy realisation.
The perspective for physics beyond the Standard Model (BSM) is reinforced by tensions emerged between some SM predictions and measurements. At present they do not reach the needed significance to claim the failure of the theory, but all together they could hint new physics (NP). Some tensions show up in the flavour sector (the flavour anomalies), identifying flavour physics as a path to disclose the BSM dynamics. Indeed, in flavour processes energy scales much higher that those directly accessed at colliders can be probed, in modes where heavy particles contribute as virtual states. To fully exploit this possibility, progress is required in the theoretical precision, as well as the increase in luminosity of the experimental facilities.
The flavour changing neutral current (FCNC) modes are of particular interest. In the SM they are forbidden at tree-level, as a consequence of the universality of the weak interactions and of the unitarity of the Cabibbo-Kobayashi-Maskawa matrix (CKM) rotating quark flavour eigenstates to mass eigenstates.  [^1] A complete discussion is in [2]. Among the tensions, there are indeed anomalies in observables related to the FCNC $b\to s$ transition. Moreover, there are difficulties concerning the value of some CKM matrix elements. Hints of violation of lepton flavour universality have also been spotted. A more detailed list of discrepancies includes [3, 4]:
  • Tensions related to the CKM matrix elements $|V_{cb}|$ and $|V_{ub}|$. In both cases the value measured from exclusive decay modes (namely $B\to D^{*}\mu\bar{\nu}_{\mu}$ and $B\to\pi\mu\bar{\nu}_{\mu}$) is smaller than the value obtained from the inclusive $B\to X_{c,u}\mu\bar{\nu}_{\mu}$ modes [5]. This long standing issue is still unsolved despite the numerous proposed explanations. [^2] For $|V_{cb}|$ a solution based on $\mu-e$ universality violation has been proposed in [6]. In addition, there is the so-called Cabibbo-angle anomaly, a $\sim 3\sigma$ tension among the determinations of $|V_{us}|$, together with a deficit from unitarity in the first row of the CKM matrix [7].
  • Tensions in observables related to the $b\to s$ transition. The branching ratios ${\cal B}(B\to K\mu^{+}\mu^{-})$ and ${\cal B}(B_{s}\to\phi\mu^{+}\mu^{-})$ are below the SM estimates. For the mode $B\to K^{*}\mu^{+}\mu^{-}$ a number of angular observables can be defined, which depend on $q^{2}$, the squared dilepton invariant mass [8]. For low $q^{2}$ some observables exceed the SM prediction by about $3\sigma$ [5].
  • Anomalies in semileptonic decays induced by $b\to c\ell\bar{\nu}_{\ell}$ transition ($\ell=e,\,\mu,\,\tau$). The measured ratios of branching fractions $R(D^{(*)})={\cal B}(B\to D^{(*)}\tau\bar{\nu}_{\tau})/{\cal B}(B\to D^{(*)}\ell\bar{\nu}_{\ell})$, with $\ell=e,\,\mu$, exceed the SM result: The current average of BaBar, Belle and LHCb Collaborations measurements deviates by about $3\sigma$ from the SM prediction [5]. This issue challenges the SM property of lepton flavour universality.
In addition to the above tensions, a discrepancy appeared in $(g-2)_{\mu}$, the anomalous magnetic moment of the muon, in particular with the Muon $g-2$ Collaboration measurements at Fermilab [9]. In this observable the main error in the SM prediction is related to one of the hadronic contributions, the hadronic-vacuum-polarisation contribution, a nonperturbative QCD effect. A lattice QCD calculation of this contribution by the BMW Collaboration [10] differs by about $2.2\sigma$ from the one obtained by a data-driven dispersive method employing the hadronic $e^{+}e^{-}$ cross section, used for the SM prediction in the White Paper [11]. In this respect, the measurement of $\sigma(e^{+}e^{-}\to\pi^{+}\pi^{-})$ by the CMD-3 experiment is also in tension with previous results [12]. An updated analysis accounting for both effects shows that the discrepancy is drastically reduced [13]. Nevertheless, non SM contributions to $(g-2)_{\mu}$ deserve to be scrutinized.
A single above mentioned deviation is not large enough to unambiguously signal NP. In general, the precision analysis is a complex task, since most observables involve SM parameters and nonperturbative quantities whose uncertainty need to be reliably assessed, as demonstrated by $(g-2)_{\mu}$. Methods have been proposed to get rid, to a large extent, of the errors of some SM inputs in a class of rare $K$ and $B$ decays [14, 15]. Taken all together the tensions could hint a BSM scenario, which however is severely constrained by the unreasonable effectiveness of the Standard Model.
A minimal extension of SM has been investigated in [1], based on the additional $U(1)^{\prime}$ gauge group and comprising right-handed neutrinos. The focus was on a solution found for the set of equations enabling the cancellation of the gauge anomalies induced by the new $U(1)^{\prime}$. In this solution the $U(1)^{\prime}$ charges of the fermions are generation dependent, and are expressed in terms of the SM hypercharges and of two additional rational parameters. One consequence of this assignment is a set of correlations arising among quark observables, among lepton observables and, remarkably, in the quark-lepton sectors together. Some correlations have been investigated in [1]. The conclusion was that, within the model, quark and lepton sectors mutually act to prevent large deviations from SM, justifying the small discrepancy with SM in observables where the tensions have shown up. However, the path to new physics is still viable: it relies on lepton flavour violating processes forbidden in SM, the observation of which would be a smoking gun for BSM physics. Such a kind of processes are the main topic of the present study. We show that, within the model (from now on denoted as the ABCD model) some charged lepton flavour violating modes can occur at non negligible rates. Moreover, correlations can be established, for example among some rare exclusive decays induced by the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition and some lepton decays.  [^3] For recent analyses of lepton flavour violation in $b\to s$ modes see [16, 17, 18, 19].
The plan of the paper is as follows: In Sec. 2 we describe the main features of the ABCD model, summarising the discussion in [1]. In Sec. 3 we select the set of observables to study within the model, and identify the correlations among them to investigate. The results of the numerical analysis are presented in Sec. 4. The conclusions are in the last section. Some details are included in the appendices. Appendix A collects the definition of the $B\to K^{*}$ matrix elements of the relevant local operators in terms of form factors. Appendix B contains the expressions of the angular coefficient functions of the $B\to K^{*}(K\pi)\ell^{-}_{1}\ell^{+}_{2}$ fully differential decay distribution. Appendix C includes the parametrization and the values of the CKM and the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrices used in this study.
2 The ABCD model
The ABCD model is a minimal extension of the SM gauge group with an additional $\text{U}(1)^{\prime}$ gauge group [1]. A new neutral gauge boson $Z^{\prime}$ is predicted, with gauge coupling $g_{Z^{\prime}}$. The charge associated to the $\text{U}(1)^{\prime}$ symmetry is named $z$-hypercharge. Many NP models are based on similar simple extensions, each one with specific $z$-hypercharge assignments [20, 21, 22, 23]. From the experimental side, the assumption on the hypercharges determines the exclusion plots in the plane of the $Z^{\prime}$ production cross section versus the mass $M_{Z^{\prime}}$.   [^4] See, e.g., the review: B.A. Dobrescu and S. Willocq, ”$Z^{\prime}$-boson searches”, in [24].
$M_{Z^{\prime}}$ sets the NP scale, and it is assumed to be acquired after spontaneous breaking of the new symmetry. In the model it is not specified how the SSB happens, the only assumption is that it occurs at a much higher scale than the SM Higgs vacuum expectation value. The mixing with other neutral gauge bosons is neglected. [^5] At tree-level the mixing vanishes in models where the SM Higgs is a singlet under $\text{U}(1)^{\prime}$. Besides $Z^{\prime}$, the model comprises heavy right-handed neutrinos considered as Dirac particles.
The quark sector is the same as in the SM: $q_{L}^{i}$ are $\text{SU}(2)_{L}$ left-handed quark doublets, $u_{R}^{i},\,d_{R}^{i}$ the right-handed singlets, $i=\{1,2,3\}$ is a generation index. In the lepton sector, $\ell_{L}^{i}$ denote the left-handed lepton doublets, $\nu_{R}^{i}$ and $\,e_{R}^{i}$ the right-handed singlets. Before the electroweak SSB the $Z^{\prime}$ couplings to fermions are flavour conserving: for a generic fermion $\psi$ the interaction Lagrangian with $Z^{\prime}$ reads
${\cal L}_{\rm int}^{Z^{\prime}}(\psi)=g_{Z^{\prime}}\,z_{\psi}\,\bar{\psi}\,\gamma^{\mu}\,\psi\,Z_{\mu}^{\prime}\;,$ (1)
with $Z_{\mu}^{\prime}$ the gauge boson field, $g_{Z^{\prime}}$ the $\text{U}(1)^{\prime}$ gauge coupling and $z_{\psi}$ the $z$-hypercharge of $\psi$. Hence, in terms of left- and right-handed fields $\psi_{L(R)}$ the fermion-$Z^{\prime}$ interaction Lagrangian is
${\cal L}_{\rm int}^{Z^{\prime}}=\sum_{i,j}\big{[}(\Delta_{L}^{\psi})^{ij}\,\bar{\psi}_{L}^{i}\,\gamma^{\mu}\,\psi_{L}^{j}+(\Delta_{R}^{\psi})^{ij}\,\bar{\psi}_{R}^{i}\,\gamma^{\mu}\,\psi_{R}^{j}\big{]}\,Z_{\mu}^{\prime}\;,$ (2)
where
$(\Delta_{L(R)}^{\psi})^{ij}=g_{Z^{\prime}}\,z_{\psi_{L(R)}}\,\delta^{ij}\;.$ (3)
After the enlargement of the gauge group, the SM remains free of gauge anomalies imposing anomaly cancellation conditions. The $z$-hypercharges are required to satisfy a set of six anomaly cancellation equations (ACE) [25], suitably written defining
$z_{q}=\sum_{i=1}^{3}z_{q_{L}^{i}}\,\,,\qquad z_{u}=\sum_{i=1}^{3}z_{u_{R}^{i}}\,\,,\qquad z_{d}=\sum_{i=1}^{3}z_{d_{R}^{i}}\,\,,$ (4)
$z_{\ell}=\sum_{i=1}^{3}z_{\ell_{L}^{i}}\,\,,\qquad z_{\nu}=\sum_{i=1}^{3}z_{\nu_{R}^{i}}\,\,,\qquad z_{e}=\sum_{i=1}^{3}z_{e_{R}^{i}}\,\,,$ (5)
and
$z^{(2)}_{f}=\sum_{i=1,2,3}z^{2}_{f^{i}}\,\,,\qquad z^{(3)}_{f}=\sum_{i=1,2,3}z^{3}_{f^{i}}\,\,.$ (6)
$f^{i}$ are the fermions $f^{i}=q_{L}^{i},\,u_{R}^{i},\,d_{R}^{i},\,\ell_{L}^{i},\,e_{R}^{i},\,\nu_{R}^{i}$ for the generations $i=1,2,3$. In a notation where groups indicate the gauge bosons in triangular graphs, the anomaly cancellation equations read:
  1. 1.
    $[SU(3)_{C}]^{2}\text{U(1)}^{\prime}$:
    $A_{33z}=2z_{q}-z_{u}-z_{d}=0\,\,;$ (7)
  2. 2.
    $[SU(2)]^{2}\text{U(1)}^{\prime}$:
    $A_{22z}=3z_{q}+z_{l}=0\,\,;$ (8)
  3. 3.
[…]
    $A_{11z}=\frac{1}{6}z_{q}-\frac{4}{3}z_{u}-\frac{1}{3}z_{d}+\frac{1}{2}z_{l}-z_{e}=0\,\,;$ (9)
  4. 4.
    triangular graphs involving two gravitons and $Z^{\prime}$:
    $A_{GGz}=3[2z_{q}-z_{u}-z_{d}]+2z_{l}-z_{e}-z_{\nu}=2z_{l}-z_{e}-z_{\nu}=0\,\,,$ (10)
    (the second equation is obtained after imposing (7));
  5. 5.
[…]
    $A_{1zz}=[z^{(2)}_{q}-2z^{(2)}_{u}+z^{(2)}_{d}]-[z^{(2)}_{l}-z^{(2)}_{e}]=0\,\,;$ (11)
  6. 6.
[…]
    $A_{zzz}=3[2z^{(3)}_{q}-z^{(3)}_{u}-z^{(3)}_{d}]+[2z^{(3)}_{l}-z^{(3)}_{\nu}-z^{(3)}_{e}]=0\,\,.$ (12)
Several studies have been devoted to the rational solutions of Eqs. (7)-(12) [26]. The solution proposed in the ABCD model consists in writing the $z$-hypercharge of each fermion $f_{i}$ in terms of the SM hypercharge $y_{f_{i}}=y_{f}$ for all $i=1,2,3$ (hence generation universal) plus a rational parameter $\epsilon_{i}$ which depends on the generation $i$ and is the same for all fermions in the generation:
$z_{f_{i}}=y_{f}+\epsilon_{i}\,.$ (13)
This choice implies
$z_{f}=\sum_{i=1}^{3}z_{f_{i}}=3y_{f}+\epsilon\,\,,$ (14)
with
$\epsilon=\sum_{i=1}^{3}\epsilon_{i}\,.$ (15)
Since the SM hypercharges $y_{f}$ solve the anomaly equations, the condition for $\epsilon_{i}$ to solve Eqs. (7)-(12) is simply [1]
$\epsilon=\sum_{i=1}^{3}\epsilon_{i}=0\,\,.$ (16)
At this stage the ABCD model has four new parameters: $M_{Z^{\prime}},\,g_{Z}$, $\epsilon_{1}$ and $\epsilon_{2}$ (using Eq. (16) to get rid of $\epsilon_{3}$). Other parameters are involved in the fermion rotation from flavour to mass eigenstates. We denote by ${\hat{V}}_{L(R)}^{\psi}$ the matrix rotating left (right)-handed components of the vector of fermions $(\psi_{L(R)}^{1},\,\psi_{L(R)}^{2}\,,\psi_{L(R)}^{3})$ with the same quantum numbers in the three generations, to obtain a diagonal mass matrix $(\hat{M}_{\psi})_{D}$:
$(\hat{M}_{\psi})_{D}=({\hat{V}}_{L}^{\psi})^{\dagger}\hat{M}_{\psi}{\hat{V}}_{R}^{\psi}\,.$ (17)
In terms of the rotated fields, Eq. (2) reads:
${\cal L}_{\rm int}^{Z^{\prime}}=\sum_{i,j}g_{Z^{\prime}}\Big{[}{\bar{\psi}}_{L}^{i}\gamma^{\mu}\,({\hat{V}})^{\dagger}_{L}{\hat{Z}}^{\psi}_{L}{\hat{V}}_{L}\,\psi_{L}^{j}+{\bar{\psi}}_{R}^{i}\gamma^{\mu}\,({\hat{V}})^{\dagger}_{R}{\hat{Z}}^{\psi}_{R}{\hat{V}}_{R}\,\psi_{R}^{j}\Big{]}\,Z^{\prime}_{\mu}\,,$ (18)
where
${\hat{Z}}^{\psi}_{L(R)}=diag(z_{\psi_{L(R)}^{1}},\,z_{\psi_{L(R)}^{2}},\,z_{\psi_{L(R)}^{3}})\,.$ (19)
The Lagrangian in Eq. (18) has the same form as in (2) defining
$(\Delta^{\psi}_{L})^{ij}=g_{Z^{\prime}}[(V_{L}^{\psi})^{\dagger}\hat{Z}^{\psi}_{L}V_{L}^{\psi}]^{ij}\,\,,\quad(\Delta^{\psi}_{R})^{ij}=g_{Z^{\prime}}[(V_{R}^{\psi})^{\dagger}\hat{Z}^{\psi}_{R}V_{R}^{\psi}]^{ij}\,.$ (20)
In SM the product $({\hat{V}}_{L}^{u})^{\dagger}{\hat{V}}_{L}^{d}$ defines the CKM matrix: $({\hat{V}}_{L}^{u})^{\dagger}{\hat{V}}_{L}^{d}={\hat{V}}_{CKM}$, and it is possible to choose ${\hat{V}}_{L}^{u}=1$ and ${\hat{V}}_{L}^{d}={\hat{V}}_{CKM}$. The same choice is made in the ABCD model. The absence of right-handed charged currents and the flavour universality of the neutral currents allows also to choose ${\hat{V}}_{R}^{u}=1$ in SM. This choice is not automatically justified in the ABCD model, due to the flavour nonuniversality of the $Z^{\prime}$ mediated current: the choice ${\hat{V}}_{R}^{u}=1$ is retained as a model assumption. Hence, the quark sector only involves the matrices ${\hat{V}}_{CKM}$ and ${\hat{V}}_{R}^{d}$.
In the lepton sector the matrix ${\hat{V}}_{R}^{\nu}$ can be ignored since no phenomenological observable involving $\nu_{R}$ is considered. In analogy to the quark sector, the matrix ${\hat{V}}_{L}^{\nu}$ is set to the unit matrix and $({\hat{V}}_{L}^{\ell})^{\dagger}$ defines the PMNS matrix: $({\hat{V}}_{L}^{\ell})^{\dagger}={\hat{U}}_{PMNS}$. Hence, the remaining matrices in this sector are ${\hat{U}}_{PMNS}$ and ${\hat{V}}_{R}^{e}$.
The parametrization of the rotation matrices can be found in [1]. Appendix C contains the choice for the parameters of the CKM and PMNS matrices. After the fermion rotations, flavour violating $Z^{\prime}$ couplings are generated, as indicated below.
  • Couplings of left-handed down quarks to $Z^{\prime}$:
    $\Delta_{L}^{ij}$ $=$ $g_{Z^{\prime}}\big{[}(2\epsilon_{1}+\epsilon_{2})V_{uj}V^{*}_{ui}+(\epsilon_{1}+2\epsilon_{2})V_{cj}V^{*}_{ci}\big{]}\qquad\qquad(i\neq j)\,\,,$ (21)
    $\Delta_{L}^{ii}$ $=$ $g_{Z^{\prime}}\Big{[}\frac{1}{6}+(2\epsilon_{1}+\epsilon_{2})|V_{ui}|^{2}+(\epsilon_{1}+2\epsilon_{2})|V_{ci}|^{2}-\epsilon_{1}-\epsilon_{2}\Big{]}\,.$ (22)
    The generation index $i=d,\,s,\,b$ is used to make the notation more transparent. $V_{ui}$ and $V_{ci}$ are CKM matrix elements.
  • Couplings of right-handed down quarks to $Z^{\prime}$:
    $\Delta_{R}^{sd}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{K}}\tilde{s}_{12}\,(\epsilon_{1}-\epsilon_{2})\,,$ (23)
    $\Delta_{R}^{bd}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{d}}\tilde{s}_{13}\,(2\epsilon_{1}+\epsilon_{2})\,,$ (24)
    $\Delta_{R}^{bs}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{s}}\tilde{s}_{23}\,(\epsilon_{1}+2\epsilon_{2})\,,$ (25)
    $\Delta_{R}^{ii}$ $=$ $g_{Z^{\prime}}[-\frac{1}{3}+\epsilon_{i}]\,.$ (26)
    The real numbers ${\tilde{s}}_{12},\,{\tilde{s}}_{13},\,{\tilde{s}}_{23}$ and the phases $\phi_{K},\,\phi_{d},\,\phi_{s}$ appear in the parametrization of the rotation matrices [1]. $(\tilde{s}_{12},\,\phi_{K})$ only enter in the kaon sector, $(\tilde{s}_{13},\,\phi_{d})$ only in the $B_{d}$ sector, $(\tilde{s}_{23},\,\phi_{s})$ only in the $B_{s}$ sector.
  • Couplings of up-type quarks to $Z^{\prime}$:
    $\Delta_{L}^{ii}=g_{Z^{\prime}}[\frac{1}{6}+\epsilon_{i}]\,\,,\qquad\Delta_{R}^{ii}=g_{Z^{\prime}}[\frac{2}{3}+\epsilon_{i}]\,\,,\qquad\qquad(i=u,c,t)\,.$ (27)
  • Couplings of left-handed charged leptons to $Z^{\prime}$:
    $\Delta_{L}^{ij}$ $=$ $g_{Z^{\prime}}\,[(2\epsilon_{1}+\epsilon_{2})U_{i1}U^{*}_{j1}+(\epsilon_{1}+2\epsilon_{2})U_{i2}U^{*}_{j2}]\qquad(i\neq j)\,\,,$ (28)
    $\Delta_{L}^{ii}$ $=$ $g_{Z^{\prime}}\,[-\frac{1}{2}+(2\epsilon_{1}+\epsilon_{2})|U_{i1}|^{2}+(\epsilon_{1}+2\epsilon_{2})|U_{i2}|^{2}-\epsilon_{1}-\epsilon_{2}]\,,$ (29)
    where $i=e,\,\mu,\,\tau$, and $U_{ij}$ are elements of the PMNS matrix. Lepton flavour violating couplings arise for non-vanishing $\epsilon_{1,2}$.
  • Couplings of right-handed charged leptons to $Z^{\prime}$:
    $\Delta_{R}^{\mu e}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{12}}\tilde{t}_{12}\,(\epsilon_{1}-\epsilon_{2})\,,$ (30)
    $\Delta_{R}^{\tau e}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{13}}\tilde{t}_{13}\,(2\epsilon_{1}+\epsilon_{2})\,,$ (31)
    $\Delta_{R}^{\tau\mu}$ $=$ $g_{Z^{\prime}}\,e^{i\phi_{23}}\tilde{t}_{23}\,(\epsilon_{1}+2\epsilon_{2})\,,$ (32)
    $\Delta_{R}^{ii}$ $=$ $g_{Z^{\prime}}\,[-1+\epsilon_{i}]\,.$ (33)
    The real numbers $\tilde{t}_{ij}$ and the phases $\phi_{ij}$ are the analogous of $\tilde{s}_{ij}$ and $\phi_{K,d,s}$ in the quark sector: $(\tilde{t}_{12},\,\phi_{12})$ enter in $\mu\to e$ transitions, $(\tilde{t}_{13},\,\phi_{13})$ in $\tau\to e$ transitions, $(\tilde{t}_{23},\,\phi_{23})$ in $\tau\to\mu$ transitions. Also in this case $\epsilon_{1,2}\neq 0$ produces lepton flavour violating $Z^{\prime}$ interactions.
  • Neutrino couplings to $Z^{\prime}$:
    $\Delta_{L}^{\nu_{i}\bar{\nu}_{i}}=g_{Z^{\prime}}\,\Big{[}-\frac{1}{2}+\epsilon_{i}\Big{]}\qquad\qquad(i=\nu_{1},\nu_{2},\nu_{3})\,.$ (34)
Hence, after the fermion rotation from flavour to mass eigenstates, the free parameters in the model comprise
$g_{Z^{\prime}}\,\,,\qquad M_{Z^{\prime}}\,\,,\qquad\epsilon_{1}\,\,,\qquad\epsilon_{2}\,\,,$ (35)
with $\epsilon_{1,2}$ rational numbers, and the mixing angles and phases in the $Z^{\prime}$ couplings to right-handed fermions:
$(\tilde{s}_{12},\,\phi_{K})\,,\qquad(\tilde{s}_{13},\,\phi_{d})\,,\qquad(\tilde{s}_{23},\,\phi_{s})\,$ (36)
for quarks, and
$(\tilde{t}_{12},\,\phi_{12})\,,\qquad(\tilde{t}_{13},\,\phi_{13})\,,\qquad(\tilde{t}_{23},\,\phi_{23})$ (37)
for leptons. The $Z^{\prime}$ couplings to left-handed fermions involve the parameters of the CKM and PMNS matrices, respectively in the quark and lepton sectors.
Since the number of parameters, even in this minimal extension of the Standard Model, is not small, a strategy for their treatment needs to be envisaged.
The general case where all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in Eqs.(36),(37) are taken to be different from zero was considered, allowing for $Z^{\prime}$ couplings to all right-handed fermions [1]. This possibility, denoted as the scenario C, appeared not efficiently constrained. 666 In particular, the efficient constraints imposed by the $\Delta F=2$ observables discussed in the following were not imposed. Another strategy was adopted in two other cases. First, all $\tilde{t}_{ij}$ in (37) were assumed to vanish, enabling flavour violating couplings of $Z^{\prime}$ to right-handed fermions only for quarks, a case denoted as scenario B. Finally, all $\tilde{s}_{ij}$ and $\tilde{t}_{ij}$ in (36),(37) were set to zero, enabling no flavour violating couplings of $Z^{\prime}$ to right-handed fermions [1]. In this last case, the scenario A, the new parameters are only those in (35) together with the elements of the CKM and PMNS matrices. This is the most economic scenario assumed in our study. In the following, we select a set of relevant observables, we discuss how the parameter space is constrained and the implications in a few lepton flavour conserving (LFC) and lepton flavour violating processes (LFV). Since all $Z^{\prime}$ couplings to fermions involve the two parameters $\epsilon_{1,2}$, correlations among different modes can be established: we investigate some of them, in particular the correlations between quark and lepton observables.
3 Observables
The couplings (21)-(34) show that in the ABCD model flavour changing neutral current transitions occur at tree-level through $Z^{\prime}$ exchanges, and that lepton flavour violating processes are possible. For this reason, we study the $b\to s\ell_{1}^{-}\ell_{2}^{+}$ transition, both for $\ell_{1}=\ell_{2}$ and $\ell_{1}\neq\ell_{2}$, focusing on the exclusive modes $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$, $B\to K^{*}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$ for which a wealth of experimental information is available and some tensions are spotted. Moreover, we describe the leptonic LFV modes $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, and the $Z^{\prime}$ contribution to $(g-2)_{\mu}$.
3.1 Effective Hamiltonian for $b\to s\ell_{1}^{-}\ell_{2}^{+}$
The low-energy $\Delta B=-1$, $\Delta S=1$ Hamiltonian governing $b\to s\ell_{1}^{-}\ell_{2}^{+}$, comprising the SM operators for $\ell_{1}=\ell_{2}$, and operators that can be generated in BSM models with $Z^{\prime}$, has the form
[…]
$G_{F}$ is the Fermi constant, $V_{ij}$ are elements of the CKM matrix, and doubly Cabibbo suppressed terms proportional to $V_{ub}V_{us}^{*}$ are neglected. [^7] The complete basis of operators including (pseudo-)scalar and tensor operators can be found in [8]. The effective Hamiltonian includes the current-current operators $O_{1,2}$
$O_{1}=({\bar{c}}_{\alpha}\gamma_{\mu}P_{L}\,b_{\beta})({\bar{s}}_{\beta}\gamma^{\mu}P_{L}\,c_{\alpha})\,\,,\quad O_{2}=({\bar{c}}\gamma_{\mu}P_{L}\,b)({\bar{s}}\gamma^{\mu}P_{L}\,c)\,\,,$ (39)
the QCD penguin operators $O_{3,\dots,6}$
$O_{3}$ $=$ $({\bar{s}}\gamma^{\mu}P_{L}\,b)\sum_{q}({\bar{q}}\gamma^{\mu}P_{L}\,q)\,\,,\quad O_{4}=({\bar{s}}_{\alpha}\gamma^{\mu}P_{L}\,b_{\beta})\sum_{q}({\bar{q}}_{\beta}\gamma^{\mu}P_{L}\,q_{\alpha})\,\,,$ (40)
$O_{5}$ $=$ $({\bar{s}}\gamma^{\mu}P_{L}\,b)\sum_{q}({\bar{q}}\gamma^{\mu}P_{R}\,q)\,\,,\quad O_{6}=({\bar{s}}_{\alpha}\gamma^{\mu}P_{L}\,b_{\beta})\sum_{q}({\bar{q}}_{\beta}\gamma^{\mu}P_{R}\,q_{\alpha})\,\,,$ (41)
with $P_{R,L}=\displaystyle\frac{1\pm\gamma_{5}}{2}$, $\alpha$,$\beta$ colour indices, and the sum in (40)-(41) running over the flavours $q=u,d,s,c,b$. The other operators in (38) are the magnetic penguin operators
$O_{7}$ $=$ $\frac{e}{16\pi^{2}}\Big{(}{\bar{s}}\sigma^{\mu\nu}(m_{s}P_{L}+m_{b}P_{R})\,b\Big{)}F_{\mu\nu}\,\,,$ (42)
$O_{7}^{\prime}$ $=$ $\frac{e}{16\pi^{2}}\Big{(}{\bar{s}}\sigma^{\mu\nu}(m_{s}P_{R}+m_{b}P_{L})\,b\Big{)}F_{\mu\nu}\,\,,$ (43)
$O_{8}$ $=$ $\frac{g_{s}}{16\pi^{2}}\Bigg{(}{\bar{s}}_{\alpha}\sigma^{\mu\nu}\Big{(}{\lambda^{a}\over 2}\Big{)}_{\alpha\beta}(m_{s}P_{L}+m_{b}P_{R})b_{\beta}\Bigg{)}G^{a}_{\mu\nu}\,\,,$ (44)
$O_{8}^{\prime}$ $=$ $\frac{g_{s}}{16\pi^{2}}\Bigg{(}{\bar{s}}_{\alpha}\sigma^{\mu\nu}\Big{(}{\lambda^{a}\over 2}\Big{)}_{\alpha\beta}(m_{s}P_{R}+m_{b}P_{L})b_{\beta}\Bigg{)}G^{a}_{\mu\nu}\,\,,$ (45)
and the semileptonic electroweak penguin operators
$O_{9}$ $=$ ${e^{2}\over 16\pi^{2}}\Big{(}{\bar{s}}\gamma^{\mu}P_{L}\,b\Big{)}\;{\bar{\ell}}_{1}\gamma_{\mu}\ell_{2}\,\,,\hskip 39.83368ptO_{9}^{\prime}={e^{2}\over 16\pi^{2}}\Big{(}{\bar{s}}\gamma^{\mu}P_{R}\,b\Big{)}\;{\bar{\ell}}_{1}\gamma_{\mu}\ell_{2}\,\,,$ (46)
$O_{10}$ $=$ ${e^{2}\over 16\pi^{2}}\Big{(}{\bar{s}}\gamma^{\mu}P_{L}\,b\Big{)}\;{\bar{\ell}}_{1}\gamma_{\mu}\gamma_{5}\ell_{2}\,\,,\hskip 28.45274ptO_{10}^{\prime}={e^{2}\over 16\pi^{2}}\Big{(}{\bar{s}}\gamma^{\mu}P_{R}\,b\Big{)}\;{\bar{\ell}}_{1}\gamma_{\mu}\gamma_{5}\ell_{2}\,\,.$ (47)
Primed operators have opposite chirality with respect to the unprimed ones. $\lambda^{a}$ are the Gell-Mann matrices, $F_{\mu\nu}$ and $G^{a}_{\mu\nu}$ the electromagnetic and gluonic field strengths, $e$ and $g_{s}$ the electromagnetic and strong coupling constant. The $b$ and $s$ quark mass $m_{b}$ and $m_{s}$ are defined in the ${\rm\overline{MS}}$ scheme at the scale $m_{b}$.
In SM, at the leading order in $\alpha_{s}$ the only operators to be considered are $O_{7}$ and $O_{9,10}$ with $\ell_{1}=\ell_{2}$. At this order, $O_{7}$ is the only operator contributing to $b\to s\gamma$. The renormalization group evolution to the scale $\mu_{b}\simeq{\cal O}(m_{b})$ also involves the magnetic gluon penguin operator $O_{8}$ and the operators $O_{1,\dots 6}$. Their mixing into $O_{7}$ generates large logarithms producing a strong enhancement of the rates. The anomalous dimension matrix governing the mixing is regularization scheme dependent. One can get rid of such a dependence defining an effective coefficient $C_{7}^{(0)\rm eff}(\mu_{b})$ which includes contributions of $O_{1,\dots 6}$ [27]. Non factorizable contributions to the matrix elements of $O_{1}$ and $O_{2}$ must also be taken into account. Their effects produce a further shift in the coefficient of $O_{7}$: $C_{7}^{\rm eff}=1.33\,C_{7}^{(0)\rm eff}$ [28]. In this way, $O_{7}$ turns out to give the dominant contribution to $b\to s\gamma$, with the SM Wilson coefficients known at NNLO in QCD [29, 2]. $O_{7}$ contributes to $b\to s\ell^{+}_{1}\ell^{-}_{2}$ for $\ell_{1}=\ell_{2}$, due to the photon conversion into the lepton pair.
In SM, the contribution to $b\to s\ell^{+}\ell^{-}$ comes from the operators $O_{9}$ and $O_{10}$ and also from other operators in (38), in particular $O_{1}$ and $O_{2}$ due to charm re-scattering. This contribution overwhelms the contributions of the other operators in the region of $q^{2}$, the dilepton invariant mass, corresponding to the narrow charmonium resonances $J/\psi$, $\psi(2S)$, …The contribution of charm re-scattering in the other intervals of $q^{2}$, the role of which has been pointed out long ago [30], has theoretical uncertainties difficult to evaluate. A possibility to account for it is to replace $C_{9}$ with $C_{9}^{\rm eff}=C_{9}+Y(q^{2})$ [31, 32, 33, 34, 35], introducing a function $Y(q^{2})$ which comprises the terms described in [36]. This could obscure genuine new short-distance contributions to $C_{9}$. Charm re-scattering contributions can be described through non-local form factors determined, e.g., by dispersion relations employing information from the nonleptonic modes $B\to K^{*}(J/\psi,\psi(2S))$ [37, 38, 39, 36], with an uncertainty difficult to estimate [40]. In the comparison with measurements it is customary to exclude the $q^{2}$ region corresponding to the resonances $J/\psi$, $\psi(2S)$, which are cut in the experimental analyses to isolate the short-distance part of the $q^{2}$ spectrum [41]. Since charm re-scattering does not affect the LFV processes with $\ell_{1}\neq\ell_{2}$, of prime interest in our study, for the $\ell^{+}\ell^{-}$ mode we also cut the resonance region and omit to include charm rescattering effects in the remaining $q^{2}$ interval.
Summarizing, we only consider the operators $O_{7}$, $O_{9,10}$ and $O^{\prime}_{9,10}$ in the effective Hamiltonian (38) to analyse the exclusive modes $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$, $B\to K^{*}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$. In SM the Wilson coefficients are flavour universal, in the ABCD extension they depend on the lepton flavour.
3.1.1 $B_{s}\to\ell_{1}^{-}\ell_{2}^{+}$
For the decay $B_{s}(p)\to\ell_{1}^{-}(k_{1})\ell_{2}^{+}(k_{2})$, with $p,k_{1,2}$ the particle four-momenta, a single hadronic quantity is required, the decay constant $f_{B_{s}}$ defined by
$\langle 0|{\bar{s}}\gamma_{\mu}\gamma_{5}b|B_{s}(p)\rangle=i\,f_{B_{s}}p_{\mu}\,\,.$ (48)
The branching fraction is given by
¯${\bar{\cal B}}(B_{s}\to\ell^{-}_{1}\ell_{2}^{+})$ $=$ $\frac{\tau_{B_{s}}}{(1-y_{s})}\frac{G_{F}^{2}\alpha^{2}|\lambda_{ts}^{*}|^{2}f_{B_{s}}^{2}}{64\pi^{3}m_{B_{s}}^{3}}\lambda^{1/2}(m_{B_{s}}^{2},\,m_{\ell_{1}}^{2},\,m_{\ell_{2}}^{2})$ (49)
$\times$ $\Big{\{}|C_{9}-C_{9}^{\prime}|^{2}(m_{\ell_{1}}-m_{\ell_{2}})^{2}\,[m_{B_{s}}^{2}-(m_{\ell_{1}}+m_{\ell_{2}})^{2}]$
$+$ $|C_{10}-C_{10}^{\prime}|^{2}(m_{\ell_{1}}+m_{\ell_{2}})^{2}\,[m_{B_{s}}^{2}-(m_{\ell_{1}}-m_{\ell_{2}})^{2}]\Big{\}}\,\,,$
with $\lambda_{ts}^{*}=V_{tb}V_{ts}^{*}$, $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$ the Källen function, $\tau_{B_{s}}$ the average $B_{s}$ lifetime. The notation ${\bar{\cal B}}$ indicates that the effects of the $B_{s}-\bar{B}_{s}$ oscillations are taken into account through the factor $1/(1-y_{s})$ [42], with $y_{s}=0.061\pm 0.009$ [43]. In SM with $\ell_{1}=\ell_{2}$ Eq. (49) shows that only the operator $O_{10}$ contributes to this decay.
3.1.2 $\bar{B}^{0}\to\bar{K}^{*0}\ell_{1}^{-}\ell_{2}^{+}$ and $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$
Let us consider $\bar{B}^{0}(p)\to\bar{K}^{*0}(p^{\prime},\,\epsilon)\ell_{1}^{-}(k_{1})\ell_{2}^{+}(k_{2})$ (the description of $B_{s}\to\phi\ell_{1}^{-}\ell_{2}^{+}$ is analogous). $\epsilon$ is the $K^{*}$ polarization four-vector, $q=k_{1}+k_{2}$. The local form factors parametrising the matrix elements of the operators in (38) are defined in appendix A.
Figure 1: Kinematics of the decay $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\ell_{1}^{-}\ell_{2}^{+}$.
Considering the subsequent $K^{*}\to K\pi$ transition and the kinematics in Fig. 1, the fully differential decay distribution in $q^{2}$ and in the angles $\theta_{\ell},\,\theta_{K^{*}},\,\phi$ has the expression
$\frac{d^{4}\Gamma(\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\ell_{1}^{-}\ell_{2}^{+})}{dq^{2}\,d\cos\theta_{\ell}\,d\cos\theta_{K^{*}}\,d\phi}=\frac{9}{32\pi}I(q^{2},\theta_{\ell},\theta_{K^{*}},\phi)\,.$ (50)
The function $I(q^{2},\theta_{\ell},\theta_{K^{*}},\phi)$ can be written in terms of angular coefficient functions $I_{i}^{(a)}(q^{2})$ using the decomposition [8]
$I(q^{2},\theta_{\ell},\theta_{K^{*}},\phi)$ $=I_{1}^{s}(q^{2})\sin^{2}\theta_{K^{*}}+I_{1}^{c}(q^{2})\cos^{2}\theta_{K^{*}}+\big{[}I_{2}^{s}(q^{2})\sin^{2}\theta_{K^{*}}+I_{2}^{c}(q^{2})\cos^{2}\theta_{K^{*}}\big{]}\cos 2\theta_{\ell}$
$+I_{3}(q^{2})\sin^{2}\theta_{K^{*}}\sin^{2}\theta_{\ell}\cos 2\phi+I_{4}(q^{2})\sin 2\theta_{K^{*}}\sin 2\theta_{\ell}\cos\phi$
$+I_{5}(q^{2})\sin 2\theta_{K^{*}}\sin\theta_{\ell}\cos\phi$ (51)
$+\big{[}I_{6}^{s}(q^{2})\sin^{2}\theta_{K^{*}}+{I_{6}^{c}(q^{2})\cos^{2}\theta_{K^{*}}}\big{]}\cos\theta_{\ell}+I_{7}(q^{2})\sin 2\theta_{K^{*}}\sin\theta_{\ell}\sin\phi$
$+I_{8}(q^{2})\sin 2\theta_{K^{*}}\sin 2\theta_{\ell}\sin\phi+I_{9}(q^{2})\sin^{2}\theta_{K^{*}}\sin^{2}\theta_{\ell}\sin 2\phi\,.$
$I_{i}^{(a)}$ depend on the Wilson coefficients in (38), hence they probe the SM. They also depend on the hadronic form factors. Their expressions in terms of amplitudes for the various $K^{*}$ polarizations and lepton chiralities are in appendix B. Integration over the three angles $\phi,\theta_{\ell}$ and $\theta_{K^{*}}$ provides the $q^{2}$ spectrum:
$\frac{d\Gamma}{dq^{2}}=\frac{1}{4}\Big{[}6I_{1}^{s}(q^{2})+3I_{1}^{c}(q^{2})-2I_{2}^{s}(q^{2})-I_{2}^{c}(q^{2})\Big{]}\,\,,$ (52)
further integration over the physical range $q^{2}\in[(m_{\ell_{1}}+m_{\ell_{2}})^{2},(m_{B}-m_{K^{*}})^{2}]$ gives the decay width.
The hadronic uncertainties can obscure the sensitivity to NP, therefore it is convenient to define observables in which such uncertainties to a large extent cancel out. A possibility is to consider the distribution analogous to (3.1.2) for the CP conjugate decay mode, with angular coefficient functions ${\bar{I}}_{i}^{(a)}$, and define CP-even and CP-odd quantities:
$S_{i}^{(a)}(q^{2})$ $=$ $\left(I_{i}^{(a)}(q^{2})+{\bar{I}}_{i}^{(a)}(q^{2})\right)/\left(d\Gamma/dq^{2}+d{\bar{\Gamma}}/dq^{2}\right)\,\,,$ (53)
$A_{i}^{(a)}(q^{2})$ $=$ $\left(I_{i}^{(a)}(q^{2})-{\bar{I}}_{i}^{(a)}(q^{2})\right)/\left(d\Gamma/dq^{2}+d{\bar{\Gamma}}/dq^{2}\right)\,\,.$ (54)
$d{\bar{\Gamma}}/dq^{2}$ is the distribution analogous to (52) for the CP conjugated mode. Starting from such expressions, several observables can be built. In SM, $S_{7},\,S_{8}$ and $S_{9}$ are tiny.
We focus on the forward-backward asymmetry with respect to the direction of the emitted lepton
$A_{FB}(q^{2})=\frac{3}{8}[2S_{6}^{s}(q^{2})+S_{6}^{c}(q^{2})]\,\,,$ (55)
on the fraction of longitudinally and transversely polarized $K^{*}$
$F_{L}(q^{2})=-S_{2}^{c}(q^{2})\,\,,\qquad F_{T}(q^{2})=1-F_{L}(q^{2})\,\,,$ (56)
and on the angular $P$-observables [44]
$P_{1}(q^{2})$ $=$ $\displaystyle\frac{2}{F_{T}(q^{2})}S_{3}(q^{2})\,\,,\quad P_{2}(q^{2})=\displaystyle\frac{1}{2F_{T}(q^{2})}S_{6s}(q^{2})\,\,,\quad P_{3}(q^{2})=-\displaystyle\frac{1}{F_{T}(q^{2})}S_{9}(q^{2})\,\,,\quad$
$P^{\prime}_{4,5,6}(q^{2})$ $=$ $\displaystyle\frac{1}{\sqrt{F_{L}(q^{2})\,F_{T}(q^{2})}}S_{4,5,7}(q^{2})\,\,.$ (57)
The forward-backward asymmetry is recognised to be sensitive to NP, in particular because in SM it changes sign in the physical $q^{2}$ range [45, 46, 47, 35, 48], with the zero $q_{0}^{2}$ determined by
${\rm Re}[C_{9}]+{m_{b}\over q_{0}^{2}}{\rm Re}[C_{7}]\left[(m_{B}+m_{K^{*}}){T_{1}(q_{0}^{2})\over V(q_{0}^{2})}+(m_{B}-m_{K^{*}}){T_{2}(q_{0}^{2})\over A_{1}(q_{0}^{2})}\right]=0\,\,\,.$ (58)
A model independent prediction for $q_{0}^{2}$ is obtained in the large energy limit of $K^{*}$, due to spin symmetry relations among the form factors [49]
${T_{1}(E)\over V(E)}={m_{B}\over m_{B}+m_{K^{*}}}\,\,\,,\hskip 28.45274pt{T_{2}(E)\over A_{1}(E)}={(m_{B}+m_{K^{*}})\over m_{B}}\,,$ (59)
where $E$ is the $K^{*}$ energy in the $B$ rest frame. [^8] The relations (59) are broken by hard gluon corrections to the weak vertex and by hard spectator interactions [50]. Since $m_{K^{*}}^{2}\ll m_{B}^{2}$, Eqs. (58),(59) give
${\rm Re}\,[C_{9}]q_{0}^{2}+2m_{B}m_{b}{\rm Re}[C_{7}]=0\,.$ (60)
Deviations from this result would signal non SM operators in the effective Hamiltonian or/and modified values of the Wilson coefficients. Another observable
$A_{T}^{(2)}(q^{2})=\frac{2S_{3}(q^{2})}{1-F_{L}(q^{2})}$ (61)
is sensitive to the virtual photon polarization [51, 52], which in SM is mainly left-handed, the right-handed polarization being suppressed by $m_{s}/m_{b}$. Deviations are expected in scenarios where the photon can also be right-handed, an effect encoded in the operator $O_{7}^{\prime}$, which is not generated in the scenario A of the ABCD model.
3.2 LFV lepton decays
At present, the LFV modes $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ are under intense theoretical and experimental scrutiny. Without neutrino oscillations such processes are absent in SM. In a minimal extension allowing for neutrino masses and mixing, they can occur at rates much smaller than the present experimental upper bounds discussed below.
  • $\mu^{-}\to e^{-}\gamma$
    In a version of the SM extended with the see-saw mechanism to describe massive nutrinos this process can occur though a penguin diagram with internal $W$ and oscillating neutrinos. The photon is emitted by $W$. If there is a $Z^{\prime}$ with LFV couplings, other penguin diagrams contribute, with $Z^{\prime}$ and a charged lepton in the loop. The branching fraction is [53]
    $\mathcal{B}(\mu\to e\gamma)=\frac{3(4\pi)^{3}\alpha}{4G^{2}_{F}}\left[|A^{M}_{e\mu}|^{2}+|A^{E}_{e\mu}|^{2}\right],$ (62)
    where
    $A^{M}_{e\mu}=-\frac{1}{96\pi^{2}M^{2}_{Z^{\prime}}}\sum_{f=e,\mu,\tau}\left[\Delta^{fe*}_{V}\Delta^{f\mu}_{V}\left(1-3\frac{m_{f}}{m_{\mu}}\right)+\Delta^{fe*}_{A}\Delta^{f\mu}_{A}\left(1+3\frac{m_{f}}{m_{\mu}}\right)\right]\,\,,$ (63)
    $A^{E}_{e\mu}=\frac{i}{96\pi^{2}M^{2}_{Z^{\prime}}}\sum_{f=e,\mu,\tau}\left[\Delta^{fe*}_{A}\Delta^{f\mu}_{V}\left(1-3\frac{m_{f}}{m_{\mu}}\right)+\Delta^{fe*}_{V}\Delta^{f\mu}_{A}\left(1+3\frac{m_{f}}{m_{\mu}}\right)\right]\,\,,$ (64)
    and $\Delta^{ij}_{V}$ and $\Delta^{ij}_{A}$ defined as
    $\Delta_{V}^{ij}=\Delta_{R}^{ij}+\Delta_{L}^{ij}\,\,,\qquad\Delta_{A}^{ij}=\Delta_{R}^{ij}-\Delta_{L}^{ij}\,\,.$ (65)
    The sum in (63),(64) is over the charged leptons in the loop.
  • $\ell_{j}\to\ell_{i}\bar{\ell}_{l}\ell_{k}$
    In the see-saw extended SM these modes are also described by loop diagrams, either penguins or boxes. In the ABCD model a tree-level diagram contributes with $Z^{\prime}$ as mediator. The generated operators and their Wilson coefficients are
    $[O_{VXY}]^{ij}_{kl}=(\bar{\psi}_{i}\gamma^{\mu}P_{X}\psi_{j})(\bar{\psi}_{k}\gamma_{\mu}P_{Y}\psi_{l})\,\,,\qquad[C_{VXY}]^{ij}_{kl}=\frac{\Delta^{ij}_{X}\Delta^{kl}_{Y}}{M^{2}_{Z^{\prime}}}\,\,,$ (66)
    with $X,Y=L,R$. The branching fraction of $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$ is given by [1]:
    $\mathcal{B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})=\frac{m_{\tau}^{5}}{1536\pi^{3}\Gamma_{\tau}}\left[2|C_{VLL}|^{2}+2|C_{VRR}|^{2}+|C_{VLR}|^{2}+|C_{VRL}|^{2}\right]^{\mu\tau}_{\mu\mu}\,,$ (67)
    with $\Gamma_{\tau}$ the $\tau$ full decay width and a $1/2$ factor due to the identical leptons in the final state. Analogous expressions hold for $\tau^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}e^{+}e^{-}$.
3.3 $(g-2)_{e,\mu}$
The $Z^{\prime}$ contribution to the muon anomalous magnetic moment $a_{\mu}=\frac{(g-2)_{\mu}}{2}$ has been obtained in [1] using the general formulae in [53]:
$\Delta a_{\mu}\simeq-\frac{1}{16\pi^{2}}\frac{m_{\mu}^{2}}{M^{2}_{Z^{\prime}}}\sum_{f=e,\mu,\tau}\left[\left|\Delta^{f\mu}_{V}\right|^{2}\left(\frac{2}{3}-\frac{m_{f}}{m_{\mu}}\right)+\left|\Delta^{f\mu}_{A}\right|^{2}\left(\frac{2}{3}+\frac{m_{f}}{m_{\mu}}\right)\right]\,\,.$ (68)
$\Delta^{ij}_{V(A)}$ are defined in (65), and the sum is over internal charged leptons in the loop. For the electron, $\Delta a_{e}$ is obtained replacing $\mu$ by $e$ in (68).
4 Numerical analysis, results and correlations
As mentioned in Sec. 2, in the ABCD model various scenarios have been considered for the $Z^{\prime}$ couplings to fermions [1]. One of them, the scenario A, consists in assuming flavour violating couplings of $Z^{\prime}$ only with left-handed fermions, obtained setting to zero the parameters in (36)-(37). The present study is based on this minimal scenario.
[…]
$|V_{cb}|\in\big{[}|V_{cb}|_{exc},\,|V_{cb}|_{inc}\big{]}\,\,,\qquad|V_{ub}|\in\big{[}|V_{ub}|_{exc},\,|V_{ub}|_{inc}\big{]}\,\,,$ (69)
instead of choosing some preferred value. All values are quoted in the appendix C, together with the parameters of the PMNS matrix.
Figure 2: Allowed region in the $(\epsilon_{1},\,\epsilon_{2})$ plane after imposing $\Delta F=2$ constraints, for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel).
Figure 3: Allowed region in the $|V_{cb}|,\,|V_{ub}|$ plane after imposing the $\Delta F=2$ constraints (dark blue points) for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. The light shaded area corresponds to the ranges (69).
[…]
The allowed regions in the $(\epsilon_{1},\,\epsilon_{2})$ plane are displayed in Fig. 2 for the two values of $M_{Z^{\prime}}$. The regions are broader than those found in [1], a consequence of scanning $|V_{cb}|$ and $|V_{ub}|$ in the exclusive-inclusive ranges instead of choosing a preferred value. A remarkable result, shown in Fig. 3, is that the inclusive value $|V_{ub}|_{incl}$ in (69) is excluded by the $\Delta F=2$ constraints, which only allow for a smaller value. Provided with the regions of parameters, we can analyse the above listed processes.
4.1 $b\to s\ell_{1}^{+}\ell_{2}^{-}$
In SM where only $\ell_{1}=\ell_{2}$ is allowed, the values of the Wilson coefficients of the operators $O_{7}$ and $O_{9,10}$ in the low-energy Hamiltonian (38) read [55]:
$C_{7}^{\rm eff}(m_{b})=-0.450\,\,,\hskip 28.45274ptC_{9}(m_{b})=4.273\,\,,\hskip 28.45274ptC_{10}(m_{b})=-4.166\,\,,$ (70)
with $C_{7}^{\rm eff}$ defined in Sec. 3.1. Efforts have been spent in global analyses to identify the coefficients deviating from the SM value, and to which extent, to reproduce the experimental data [56, 57, 58, 59].
Figure 4: Correlation between real and imaginary parts of the $Z^{\prime}$ contributions to the Wilson coefficients $C_{9}$ (left panels) and $C_{10}$ (right panels) for the LFC $b\to s\mu^{-}\mu^{+}$ (top panels) and LFV $b\to s\mu^{-}\tau^{+}$ transition (bottom panels). The $Z^{\prime}$ mass is set to $M_{Z^{\prime}}=1$ TeV.
Figure 5: Correlation between real and imaginary parts of the $Z^{\prime}$ contributions to the Wilson coefficients $C_{9}$ and $C_{10}$ as in Fig. 4, for $M_{Z^{\prime}}=3$ TeV.
In the scenario A of the ABCD model there are no contributions to the coefficients of the primed operators in (38). The $Z^{\prime}$ contribution to $C_{7}$ is negligible with respect to SM [54]. The contributions to $C_{9}$ and $C_{10}$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV are displayed in Figs. 4 and 5, where each point corresponds to a pair real/imaginary part of the coefficients computed for the same values of the parameters. The contributions to the real parts can be sizeable, up to ±$\pm 10\%$ of the SM values. Imaginary parts of ${\cal O}(10^{-4})$ are generated, originating from the phases of the CKM and PMNS matrices. The correlations between the real parts of the coefficients are displayed in the plots in Fig. 6 for $M_{Z^{\prime}}=1$ and $3$ TeV.
The values of $C_{9}^{NP}$ and $C_{10}^{NP}$ obtained for the lepton flavour violating modes $b\to s\mu^{-}\tau^{+}$, together with their correlation, are displayed in the same figures 4,5 and 6. The anticorrelation between $(C_{9}^{NP})_{\mu\tau}$ and $(C_{10}^{NP})_{\mu\tau}$ is a feature of the scenario A of the ABCD model. Indeed, the expressions of the coefficients are
$(C_{9}^{NP})_{\ell_{1}\ell_{2}}=-\frac{1}{g_{SM}^{2}s_{W}^{2}M_{Z^{\prime}}^{2}\lambda_{ts}^{*}}\big{(}\Delta_{L}^{bs}\big{)}^{*}\Delta_{V}^{\ell_{1}\ell_{2}}\,\,,\qquad(C_{10}^{NP})_{\ell_{1}\ell_{2}}=-\frac{1}{g_{SM}^{2}s_{W}^{2}M_{Z^{\prime}}^{2}\lambda_{ts}^{*}}\big{(}\Delta_{L}^{bs}\big{)}^{*}\Delta_{A}^{\ell_{1}\ell_{2}}\,\,,$ (71)
with $g_{SM}^{2}=\frac{4G_{F}}{\sqrt{2}}\frac{e^{2}}{8\pi^{2}s_{W}^{2}}$, $s_{W}$ the sine of the Weinberg angle, and $\Delta_{V(A)}^{\ell_{1}\ell_{2}}$ defined in (65). In the scenario A where the $Z^{\prime}$ flavour violating couplings to right-handed fermions vanish, we have $\Delta_{V}^{\ell_{1}\ell_{2}}=-\Delta_{A}^{\ell_{1}\ell_{2}}$ for $\ell_{1}\neq\ell_{2}$.
Figure 6: Correlation between the real parts of the $Z^{\prime}$ contributions to Wilson coefficients $(C_{9})_{\mu\mu}$ and $(C_{10})_{\mu\mu}$(left panels) and $(C_{9})_{\mu\tau}$ and $(C_{10})_{\mu\tau}$ (right panels)for $M_{Z^{\prime}}=1$ TeV (top panels) and for $M_{Z^{\prime}}=3$ TeV(bottom panels).
4.1.1 Exclusive lepton flavour conserving $b\to s\ell^{+}\ell^{-}$ modes
The simplest process induced by $b\to s\ell^{+}\ell^{-}$ is the purely leptonic $B_{s}$ decay, that we analyse using $f_{B_{s}}=230.3\pm 1.3$ MeV [5]. Fig. 7 shows the branching fractions ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. In SM we find ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})=(3.65\pm 0.10)\times 10^{-9}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(7.70\pm 0.10)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$. The experimental world average ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})_{\rm exp}=(3.34\pm 0.27)\times 10^{-9}$ [24] is indicated in Fig. 7. The results in the ABCD model encompass the experimental range. Moreover, requiring that ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ agrees with experiment, ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.5)\times 10^{-7}$ is predicted for $M_{Z^{\prime}}=1$ TeV, and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})=(8.2\pm 0.8)\times 10^{-7}$ for $M_{Z^{\prime}}=3$ TeV. The correlation between the two purely leptonic modes is shown in the figure, where each point corresponds to the branching fractions of $B_{s}\to\mu^{+}\mu^{-}$ and $B_{s}\to\tau^{+}\tau^{-}$ computed with the same set of parameters.
For the lepton flavour conserving modes $B\to K^{*}\ell^{+}\ell^{-}$, with $\ell=\mu,\,\tau$, we use the local form factors determined in [60], see appendix A.
Figure 7: Correlation between ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded region corresponds to the experimental world average for ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$[24].
Figure 8: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded band corresponds to the experimental result ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.4\pm 0.5)\times 10^{-7}$ [24].
Figure 8 shows the correlation between the branching fractions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})$. The vertical band corresponds to the measurement ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})_{\rm exp}=(9.4\pm 0.5)\times 10^{-7}$ [24]. In the case of $B\to K^{*}\tau^{+}\tau^{-}$ only an upper limit is known: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})<3.1\times 10^{-3}$. In the ABCD model the result for ${\cal B}(\bar{B}^{0}\to K^{*0}\mu^{+}\mu^{-})$ encompasses the experimental interval. Requiring that it lies in the measured range, the predictions ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 1.3)\times 10^{-8}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(8.6\pm 0.9)\times 10^{-8}$ are obtained for $M_{Z^{\prime}}=1$ and $3$ TeV, respectively. For the SM we have: ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})=(9.6\pm 2.0)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\tau^{-})=(9.2\pm 2.0)\times 10^{-8}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$.
[…]
Figure 9: Angular coefficient functions $\tilde{I}_{i}^{(a)}(q^{2})$ for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$. The continuous line is the SM result for the central value of the form factor parameters, the light shaded region the result in the ABCD model varying the parameters in the allowed ranges, with $M_{Z^{\prime}}=1$ TeV.
Figure 10: Angular coefficient functions for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$. The continuous line is the SM result for the central value of the form factor parameters, the light shaded region the result in the ABCD model varying the parameters in the allowed ranges, with $M_{Z^{\prime}}=1$ TeV.
For the forward-backward asymmetry and the fraction $F_{L}$ of longitudinally polarized $K^{*}$, Eqs. (55),(56), the results are displayed in Fig. 11 for $\mu$ and $\tau$. No measurement is available for $\tau$. The largest deviation from SM is found in $F_{L}$ at high $q^{2}$ in the $\tau$ mode.
Figure 11: Forward-backward asymmetry $A_{FB}(q^{2})$ (left) and $K^{*}$ polarization fraction $F_{L}(q^{2})$ (right plots) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The dark continuous line is the SM result, the shaded region the ABCD result for $M_{Z^{\prime}}=1$ TeV. The black dots in the top panels correspond to the LHCb measurement [72].
For the $P$-observables in (57) where tensions emerged [69, 70, 71, 72, 73, 74, 41], we display the most significant results in Fig. 12. The effects obtained in the ABCD model are more pronounced for $\tau$, namely for low $q^{2}$ the observable $P^{\prime}_{5}$ can differ from SM.
Figure 12: Angular $P$-observables in Eq. (57) for $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\mu^{+}\mu^{-}$ (top panels) and $\bar{B}^{0}\to\bar{K}^{*0}(K\pi)\tau^{+}\tau^{-}$ (bottom panels). The black dots in the top panels correspond to LHCb measurement [72]. Same color code as in Fig. 11.
For $B_{s}\to\phi\mu^{-}\mu^{+}$, using the form factors in [60] with parameters in Table 2, we find in SM: ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})=(11.8\pm 2.3)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$, confirming that a higher value is foreseen for this rate with respect to the measured ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}=(8.4\pm 0.4)\times 10^{-7}$ [24]. Measurements are not available for the $\tau$ mode; we find ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(1.02\pm 0.20)\times 10^{-7}\left(|\lambda_{ts}^{*}|/0.04\right)^{2}$ in SM. Fig. 13 shows the correlation between the two LFC decays in the ABCD model: the plot for $M_{Z^{\prime}}=1$ TeV shows that the range for the $\tau$ mode is ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.25)\times 10^{-7}$, for $M_{Z^{\prime}}=3$ TeV the range is ¯${\bar{\cal B}}(B_{s}\to\phi\tau^{-}\tau^{+})=(0.9\pm 0.2)\times 10^{-7}$. However, since ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{-}\mu^{+})_{\rm exp}$ is not reproduced, a puzzle remains unsolved for this mode. If the measurement is confirmed, the first issue to scrutinise is the uncertainty of the $B_{s}\to\phi$ form factors.
Figure 13: Correlation between ¯${\bar{\cal{B}}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ¯${\bar{\cal{B}}}(B_{s}\to\phi\tau^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the result ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})|_{exp}=(8.4\pm 0.4)\times 10^{-7}$ quoted in [24].
   LFV mode ×${\cal B}\times 10^{9}$ ${\cal B}^{(1)}\times 10^{9}$ ${\cal B}^{(2)}\times 10^{9}$ experiment
$B_{s}\to\tau^{+}\mu^{-}$ $0.00\div 2.10$ $0.00\div 2.10$ $0.00\div 1.60$ $<4.2\times 10^{-5}$ [24]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $0.00\div 2.90$ $0.00\div 2.90$ $0.00\div 1.15$ $<1.0\times 10^{-5}$ [24]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $0.00\div 3.43$ $0.00\div 2.60$ $<1.0\times 10^{-5}$[75]
$B_{s}\to\tau^{+}\mu^{-}$ $0.00\div 2.30$ $0.00\div 1.50$ $0.00\div 0.14$ $<4.2\times 10^{-5}$ [24]
$\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ $0.00\div 3.10$ $0.00\div 2.00$ $0.00\div 0.20$ $<1.0\times 10^{-5}$ [24]
$B_{s}\to\phi\tau^{+}\mu^{-}$ $0.00\div 3.70$ $0.00\div 0.25$ $<1.0\times 10^{-5}$[75]
Table 1: Ranges of branching fractions of LFV modes in the scenario A of the ABCD model. The first three rows correspond to $M_{Z^{\prime}}=1$ TeV, the last three rows to $M_{Z^{\prime}}=3$ TeV. ${\cal B}^{(1)}$ is obtained after imposing that the corresponding lepton flavour conserving mode lies in the experimental range; ${\cal B}^{(2)}$ is obtained after imposing also the constraint from the upper limit of ${\cal B}(\mu^{-}\to e^{-}\gamma)$. The experimental upper bounds (at $90\%$ C.L.) are quoted in the last column.
Figure 14: Correlation between the branching fractions ¯${\bar{\cal B}}(B_{s}\to\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The shaded vertical band corresponds to the experimental world average [24].
Figure 15: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ and ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). Same color code as in Fig. 8.
4.1.2 Exclusive lepton flavour violating modes $b\to s\ell_{1}^{-}\ell_{2}^{+}$ (with $\ell_{1}\neq\ell_{2}$)
A feature of the ABCD model is that correlations are generated between lepton flavour conserving and lepton flavour violating modes. In Table 1 we quote the predictions for a set of LFV modes, for $M_{Z^{\prime}}=1$ TeV and $3$ TeV. The predictions can be further constrained requiring agreement of the rates of corresponding LFC modes with experiment. In the Table we quote the unconstrained result and the one constrained by the corresponding LFC process. Visible rates for the LFV modes are possible, even though they are much smaller than the present experimental upper bounds. [^9] Recent experimental investigations of $\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-}$ are described in [76, 77]. However, the upper limits are expected to be sensibly improved in the near future.
Figure 14 shows the correlation between the branching ratios of the LFC $B_{s}\to\mu^{+}\mu^{-}$ and LFV $B_{s}\to\tau^{+}\mu^{-}$ modes for the two values of $M_{Z^{\prime}}$. The experimental upper bound for ¯${\bar{\cal B}}(B_{s}\to\tau^{+}\mu^{-})$ is quoted in Table 1. In the ABCD model the branching ratio of the LFV $\tau^{+}\mu^{-}$ mode can be comparable to that of the LFC $\mu^{+}\mu^{-}$ mode. Indeed, the decay to same flavour leptons is determined only by the $C_{10}$ term, while for $\ell_{1}\neq\ell_{2}$ also $C_{9}$ contributes. Moreover, the helicity suppression in the $\ell_{1}=\ell_{2}=\mu$ mode is lifted in the $\tau^{+}\mu^{-}$ mode producing a sizable enhancement.
While for the LFC modes the results do not sensibly change varying $M_{Z^{\prime}}$ from 1 TeV to 3 TeV, this is not true for the LFV modes. The reason is that ABCD does not predict large deviations from SM: increasing the $Z^{\prime}$ mass, the SM results are approached. For modes allowed in SM, the deviations are small even for small values of $M_{Z^{\prime}}$. On the other hand, for modes forbidden in SM, increasing $M_{Z^{\prime}}$ results in a sensible decrease of the branching fractions. The same observation holds for ${\bar{B}}^{0}\to{\bar{K}}^{*0}\tau^{+}\mu^{-}$: the branching fraction predicted in the ABCD model is quoted in Table 1 both without and with the constraint given by the measured ${\cal B}({\bar{B}}^{0}\to{\bar{K}}^{*0}\mu^{+}\mu^{-})$. A further constraint given by the upper bound of ${\cal B}(\mu^{-}\to e^{-}\gamma)$ further reduces the range for this rate. Predictions for $B_{s}\to\phi\tau^{+}\mu^{-}$ are presented in the same Table 1 and in Fig.16.
Figure 16: Correlation between ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\mu^{-})$ and ¯${\bar{\cal B}}(B_{s}\to\phi\mu^{+}\tau^{-})$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). Same color code as in Fig. 13.
4.2 Correlations among lepton flavour violating processes in the quark and lepton sectors
[…]
Figure 18: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ for $M_{Z^{\prime}}=1$ TeV (top panels) and $M_{Z^{\prime}}=3$ TeV (bottom panels). The right plots are a zoom of the left ones on the region selected imposing the experimental upper bound for ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [78]. The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees within $1\sigma$ with experiment.
The $Z^{\prime}$ couplings to leptons and to quarks involve the same couple of parameters $\epsilon_{1}$ and $\epsilon_{2}$. The consequence is that correlations can be established among observables in the different sectors. Important cases are the correlations among the LFV $b\to s\ell_{1}^{+}\ell_{2}^{-}$ modes and the LFV leptonic decays $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$.
[…]
The correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and ${\cal B}(\mu^{-}\to e^{-}\gamma)$ is displayed in Fig. 18, and similar remarks hold. In particular, the bound on ${\cal B}(\mu^{-}\to e^{-}\gamma)$ [78] reduces the allowed range for ${\cal B}(B\to K^{*}\tau^{+}\mu^{-})$, see Table 1. The same occurs for $B_{s}\to\phi\tau^{+}\mu^{-}$, as we report in the same Table.
As for the correlations involving ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$, the experimental upper bound ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})<2.1\times 10^{-8}$ [24] must be taken into account, however it does not provide significant constraints for $B_{s}$ and $B$ modes, as it can be inferred from Figs. 19. Accounting for the measurement of the rate of $B_{s}\to\mu^{+}\mu^{-}$, the largest value for ${\cal B}(\tau^{-}\to\mu^{-}\mu^{+}\mu^{-})$ is about $1.2\times 10^{-8}$, a relevant result when planning new facilities.
[…]
Figure 20: Correlation between ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ and $\Delta a_{\mu}$ for $M_{Z^{\prime}}=1$ TeV (left panel) and $M_{Z^{\prime}}=3$ TeV (right panel). The light yellow points are selected requiring that the LFC ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\mu^{+}\mu^{-})$ agrees at $1\sigma$ with experiment.
A final remark concerns the $Z^{\prime}$ contribution to the muon anomalous magnetic moment and to its correlation with other observables. The results for $\Delta a_{\mu}$ vs ${\cal B}(\bar{B}^{0}\to\bar{K}^{*0}\tau^{+}\mu^{-})$ are displayed in Fig. 20. $\Delta a_{\mu}$ turns out to be negative, and it reaches at most $\Delta a_{\mu}\sim-10^{-12}$, a size similar to subleading hadron contributions, such as the hadronic light-by-light ones [11].
5 Conclusions
The search for BSM effects is currently facing the evidence that deviations from SM predictions exist but they are small. For example, the flavour anomalies are below the required level of significance to claim the evidence of new physics. Being confident that NP exists, we have pursued the aim of justifying the absence of large deviations. We have scrutinised channels where anomalies have been detected together with processes strictly forbidden in SM, namely lepton flavour violating modes, and the correlations among them.
In the proposed simple extension of the SM, the ABCD model, the new gauge boson $Z^{\prime}$ has flavour conserving and flavour violating generation dependent couplings to fermions. The model is based on a special solution of the gauge anomaly cancellation equations, which relates the quark and lepton couplings to $Z^{\prime}$. The flavour observables in the two sectors are therefore intertwined, producing nontrivial correlations.
As noticed at the formulation of the model [1], quark and lepton sectors mutually act to prevent large deviations from SM. This is a welcome feature in view of the small discrepancies with respect to SM displayed by measurements of the flavour observables. Moreover, the possibility of having lepton flavour violation implies that modes, such as those induced by $b\to s\ell_{1}^{-}\ell_{2}^{+}$ with $\ell_{1}\neq\ell_{2}$, can occur and are correlated to LFV leptonic modes, such as $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, and to the LFC modes $b\to s\ell^{-}\ell^{+}$ as well. We summarize below our main findings and indicate the corresponding plots.
  • We have shown that the NP contribution to the Wilson coefficients ${\rm Re}(C_{9}^{\rm NP})_{\ell\ell}$ and ${\rm Re}(C_{10}^{\rm NP})_{\ell\ell}$, $\ell=\mu,\,\tau$, is $\simeq 10\%$ of the corresponding SM coefficients. The analogous LFV coefficients ${\rm Re}(C_{9}^{\rm NP})_{\mu\tau}$ and ${\rm Re}(C_{10}^{\rm NP})_{\mu\tau}$ are approximately of the same size as ${\rm Re}(C_{9(10)}^{\rm NP})_{\mu\mu}$, as shown in Figs. 4-6.
  • We have found that LFV modes could reach detectable rates, probably accessible in the near future, see Figs. 14-16.
  • The role of quark and lepton observables to mutually constrain each other is visible in the results in Table 1 and in Figs. 17-18. The experimental upper limit for ${\cal B}(\mu\to e\gamma)$ efficiently constrains the $B$ and $B_{s}$ LFV decays.
  • The branching ratio of the LFV $\tau\to 3\mu$ mode can reach values close to the present upper bound, as shown in Fig. 19.
All findings are good news for the NP searches at the facilities planned within the future strategy for particle physics.
Acknowledgements
F.D.F. thanks the coauthors of [1]. We thank A.J. Buras for very useful discussions. This study has been carried out within the INFN project (Iniziativa Specifica) SPIF. The research has been partly funded by the European Union – Next Generation EU through the research Grant No. P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Universitá e Ricerca (MUR).
Appendix A $B_{q}\to V$ local form factors
We use the standard parametrization of the $B_{q}\to V$ ($q=d,s$, $V=K^{*},\,\phi$) matrix elements of local currents in terms of form factors:
$\langle V(p^{\prime},\epsilon)|{\bar{s}}\gamma_{\mu}(1-\gamma_{5})b|B_{q}(p)\rangle$ $=$ $\epsilon_{\mu\nu\alpha\beta}\epsilon^{*\nu}p^{\alpha}p^{\prime\beta}\frac{2V(q^{2})}{m_{B_{q}}+m_{V}}$ (A.1)
$-i\left[\epsilon^{*}_{\mu}(m_{B_{q}}+m_{V})A_{1}(q^{2})-(\epsilon^{*}\cdot q)(p+p^{\prime})_{\mu}\frac{A_{2}(q^{2})}{m_{B_{q}}+m_{V}}\right.$
$-\left.(\epsilon^{*}\cdot q)\frac{2m_{V}}{q^{2}}\big{(}A_{3}(q^{2})-A_{0}(q^{2})\big{)}q_{\mu}\right]\,\,,\,\,\,\hskip 14.22636pt$
$B\to K^{*}$ $m_{R}\,({\rm GeV})$ $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$
$V$ $5.415$ $0.34\pm 0.04$ $-1.05\pm 0.24$ $2.37\pm 1.39$
$A_{1}$ $5.829$ $0.27\pm 0.03$ $0.30\pm 0.19$ $-0.11\pm 0.48$
$A_{0}$ $5.366$ $0.36\pm 0.05$ $-1.04\pm 0.27$ $1.12\pm 1.35$
$A_{12}$ $5.829$ $0.26\pm 0.03$ $0.60\pm 0.20$ $0.12\pm 0.84$
$T_{1}$ $5.415$ $0.28\pm 0.03$ $-0.89\pm 0.19$ $1.95\pm 1.10$
$T_{2}$ $5.829$ $0.28\pm 0.03$ $0.40\pm 0.18$ $0.36\pm 0.51$
$T_{23}$ $5.829$ $0.67\pm 0.08$ $1.48\pm 0.49$ $1.92\pm 1.96$
$B_{s}\to\phi$ $m_{R}\,({\rm GeV})$ $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$
$V$ $5.415$ $0.39\pm 0.03$ $-1.03\pm 0.25$ $3.50\pm 1.55$
$A_{1}$ $5.829$ $0.30\pm 0.03$ $0.48\pm 0.19$ $0.29\pm 0.65$
$A_{0}$ $5.366$ $0.39\pm 0.05$ $-0.78\pm 0.26$ $2.41\pm 1.48$
$A_{12}$ $5.829$ $0.25\pm 0.03$ $0.76\pm 0.20$ $0.71\pm 0.96$
$T_{1}$ $5.415$ $0.31\pm 0.03$ $-0.87\pm 0.19$ $2.75\pm 1.19$
$T_{2}$ $5.829$ $0.31\pm 0.03$ $0.58\pm 0.19$ $0.89\pm 0.71$
$T_{23}$ $5.829$ $0.68\pm 0.07$ $2.11\pm 0.46$ $4.94\pm 2.25$
Table 2: Parameters of the $B_{(s)}\to V$ form factors in Eq. (A.5) determined in [60].
[…]
$+\Big{[}\epsilon^{*}_{\mu}(m_{B_{q}}^{2}-m^{2}_{V})-(\epsilon^{*}\cdot q)(p+p^{\prime})_{\mu}\Big{]}\;T_{2}(q^{2})\hskip 19.91684pt$
$+(\epsilon^{*}\cdot q)\left[q_{\mu}-\frac{q^{2}}{m_{B_{q}}^{2}-m^{2}_{V}}(p+p^{\prime})_{\mu}\right]\;T_{3}(q^{2})\,,\,\,\,\,\,\,$
with $q=p-p^{\prime}$ and $\epsilon$ the $V$ polarization vector. The form factors are not all independent: $A_{3}$ is given by
$A_{3}(q^{2})=\frac{m_{B_{q}}+m_{V}}{2m_{V}}A_{1}(q^{2})-\frac{m_{B_{q}}-m_{V}}{2m_{V}}A_{2}(q^{2})$ (A.3)
with $A_{3}(0)=A_{0}(0)$. $T_{1}(0)=T_{2}(0)$ follows from the identity $\sigma_{\mu\nu}\gamma_{5}=-\frac{i}{2}\epsilon_{\mu\nu\alpha\beta}\sigma^{\alpha\beta}$ (with $\epsilon_{0123}=+1$).
In our analysis we use the determination in [60] based on light-cone QCD sum rules (see also [80]). Defining the ”conformal” variable
$z(t)=\frac{\sqrt{t_{+}-t}-\sqrt{t_{+}-t_{0}}}{\sqrt{t_{+}-t}+\sqrt{t_{+}-t_{0}}}\,$ (A.4)
with $t_{\pm}=(m_{B_{q}}\pm m_{V})^{2}$ and $t_{0}=t_{+}\left(1-\sqrt{1-\frac{t_{-}}{t_{+}}}\right)$, a generic form factor $F_{i}$ is written as the sum of the first three terms of the $z-$expansion
$F_{i}(q^{2})=\frac{1}{1-\frac{q^{2}}{m_{R,i}^{2}}}\sum_{k=0}^{2}\alpha_{k}^{i}\left[z(q^{2})-z(0)\right]^{k}\,\,,$ (A.5)
with parameters $\alpha_{k}^{i}$ and $m_{R,i}$. The value of $m_{R,i}$ is close to the mass of the lightest t-channel resonance involved in the various terms in (A.1), (A.2). The parameters are determined for $V,\,A_{1},\,A_{0},\,A_{12}$ and $T_{1},\,T_{2}$, $T_{23}$ in [60], defining
$A_{12}(q^{2})$ $=$ $\frac{(m_{B_{q}}+m_{V})^{2}(m_{B_{q}}^{2}-m_{V}^{2}-q^{2})A_{1}(q^{2})-\lambda(m_{B_{q}}^{2},m_{V}^{2},q^{2})A_{2}(q^{2})}{16m_{B_{q}}m_{V}^{2}(m_{B_{q}}+m_{V})}$
$T_{23}(q^{2})$ $=$ $\frac{(m_{B_{q}}^{2}-m_{V}^{2})(m_{B_{q}}^{2}+3m_{V}^{2}-q^{2})T_{2}(q^{2})-\lambda(m_{B_{q}}^{2},m_{V}^{2},q^{2})T_{3}(q^{2})}{8m_{B_{q}}m_{V}^{2}(m_{B_{q}}-m_{V})}\,\,.$ (A.6)
To make easier to reproduce our results, we quote the values of the parameters in Table 2, while their correlation matrix is in [60].
Appendix B Angular coefficient functions for $B\to K^{*}(K\pi)\ell_{1}^{-}\ell_{2}^{+}$
The coefficient functions $I_{i}^{(a)}$, for $i=1,\dots 9$ and $a=s,\,c$ in (3.1.2), can be written in terms of the transversity amplitudes
$A_{\perp L,R}(q^{2})$ $=$ $N(q^{2})\sqrt{2}\lambda_{H}^{1/2}\bigg{[}\left[(C_{9}^{\rm eff}+C_{9}^{{\rm eff}\prime})\mp(C_{10}^{\rm eff}+C_{10}^{{\rm eff}\prime})\right]\frac{V(q^{2})}{m_{B}+m_{K^{*}}}$
$+$ $\frac{2(m_{b}+m_{s})}{q^{2}}(C_{7}^{\rm eff}+C_{7}^{{\rm eff}\prime})T_{1}(q^{2})\bigg{]},$
$A_{\parallel L,R}(q^{2})$ $=$ $-N(q^{2})\sqrt{2}(m_{B}^{2}-m_{K^{*}}^{2})\bigg{[}\left[(C_{9}^{\rm eff}-C_{9}^{{\rm eff}\prime})\mp(C_{10}^{\rm eff}-C_{10}^{{\rm eff}\prime})\right]\frac{A_{1}(q^{2})}{m_{B}-m_{K^{*}}}$
$+$ $\frac{2(m_{b}-m_{s})}{q^{2}}(C_{7}^{\rm eff}-C_{7}^{{\rm eff}\prime})T_{2}(q^{2})\bigg{]},$
$A_{0L,R}(q^{2})$ $=$ $-\frac{N(q^{2})}{2m_{K^{*}}\sqrt{q^{2}}}\bigg{\{}\left[(C_{9}^{\rm eff}-C_{9}^{{\rm eff}\prime})\mp(C_{10}^{\rm eff}-C_{10}^{{\rm eff}\prime})\right]$
$\times$ $\bigg{[}(m_{B}^{2}-m_{K^{*}}^{2}-q^{2})(m_{B}+m_{K^{*}})A_{1}(q^{2})-\lambda_{H}\frac{A_{2}(q^{2})}{m_{B}+m_{K^{*}}}\bigg{]}$
$+$ ${2(m_{b}-m_{s})}(C_{7}^{\rm eff}-C_{7}^{{\rm eff}\prime})\bigg{[}(m_{B}^{2}+3m_{K^{*}}^{2}-q^{2})T_{2}(q^{2})-\frac{\lambda_{H}}{m_{B}^{2}-m_{K^{*}}^{2}}T_{3}(q^{2})\bigg{]}\bigg{\}},$
$A_{t\,L,R}(q^{2})$ $=$ $-\frac{N(q^{2})}{\sqrt{q^{2}}}\lambda_{H}^{1/2}\big{[}(C_{9}^{\rm eff}-C_{9}^{{\rm eff}\prime})\mp(C_{10}^{\rm eff}-C_{10}^{{\rm eff}\prime})\Big{]}A_{0}(q^{2}).$
The factor $N(q^{2})$ reads
$N(q^{2})=\lambda_{ts}^{*}\Big{[}\frac{G_{F}^{2}\alpha^{2}}{3\times 2^{10}\pi^{5}m_{B}^{3}}\lambda_{H}^{1/2}\lambda_{L}^{1/2}\Big{]}^{1/2}\,\,,$ (B.2)
with $\lambda_{H}=\lambda(q^{2},m_{B}^{2},m_{K^{*}}^{2})$ and $\lambda_{L}=\lambda(q^{2},m_{1}^{2},m_{2}^{2})$.
The expressions of the coefficient functions are:
$I_{1}^{s}$ $=$ $\frac{\lambda_{L}+2[q^{4}-(m_{1}^{2}-m_{2}^{2})^{2}]}{4q^{4}}\left[|A_{\perp L}|^{2}+|A_{\parallel L}|^{2}+(L\to R)\right]+\frac{4m_{1}m_{2}}{q^{2}}{\rm Re}\left[A_{\perp L}A_{\perp R}^{*}+A_{\parallel L}A_{\parallel R}^{*}\right],$
$I_{1}^{c}$ $=$ $\frac{q^{4}-(m_{1}^{2}-m_{2}^{2})^{2}}{q^{4}}\big{[}|A_{0L}|^{2}+|A_{0R}|^{2}\big{]}+\frac{8m_{1}m_{2}}{q^{2}}{\rm Re}\big{[}A_{0L}A_{0R}^{*}-A_{t\,L}A_{t\,R}^{*}\big{]}$
$-2\frac{(m_{1}^{2}-m_{2}^{2})^{2}-q^{2}(m_{1}^{2}+m_{2}^{2})}{q^{4}}\left[|A_{t\,L}|^{2}+|A_{t\,R}|^{2}\right],$
$I_{2}^{s}$ $=$ $\frac{\lambda_{L}}{4q^{4}}\left[|A_{\perp L}|^{2}+|A_{\parallel L}|^{2}+(L\to R)\right],$
$I_{2}^{c}$ $=$ $-\frac{\lambda_{L}}{q^{4}}\left[|A_{0L}|^{2}+(L\to R)\right],$
$I_{3}$ $=$ $\frac{\lambda_{L}}{2q^{4}}\left[|A_{\perp L}|^{2}-|A_{\parallel L}|^{2}+(L\to R)\right],$
$I_{4}$ $=$ $\frac{\lambda_{L}}{\sqrt{2}q^{4}}\left[{\rm Re}(A_{0L}A_{\parallel L}^{*})+(L\to R)\right],$ (B.3)
$I_{5}$ $=$ $\frac{\sqrt{2}\lambda_{L}^{1/2}}{q^{2}}\left[{\rm Re}[A_{0L}A_{\perp L}^{*}-(L\to R)]-\frac{m_{1}^{2}-m_{2}^{2}}{q^{2}}{\rm Re}[A_{t\,L}A_{\parallel L}^{*}+(L\to R)]\right],$
$I_{6}^{s}$ $=$ $\frac{2\lambda_{L}^{1/2}}{q^{2}}\left[{\rm Re}(A_{\parallel L}A_{\perp L}^{*})-(L\to R)\right],$
$I_{6}^{c}$ $=$ $\frac{4\lambda_{L}^{1/2}}{q^{2}}\frac{m_{1}^{2}-m_{2}^{2}}{q^{2}}\left[{\rm Re}(A_{0L}A_{t\,L}^{*})+(L\to R)\right],$
$I_{7}$ $=$ $\frac{\sqrt{2}\lambda_{L}^{1/2}}{q^{2}}\left[{\rm Im}[A_{0L}A_{\parallel L}^{*}-(L\to R)]+\frac{m_{1}^{2}-m_{2}^{2}}{q^{2}}{\rm Im}[A_{\perp L}A_{t\,L}^{*}+(L\to R)]\right],$
$I_{8}$ $=$ $\frac{\lambda_{L}}{\sqrt{2}q^{4}}\left[{\rm Im}(A_{0L}A_{\perp L}^{*})+(L\to R)\right],$
$I_{9}$ $=$ $\frac{\lambda_{L}}{q^{4}}\left[{\rm Im}(A_{\parallel L}^{*}A_{\perp L})+(L\to R)\right].$
The above expressions are modified if other BSM operators are included in the low-energy Hamiltonian, besides the ones in (38).
Appendix C Parametrization of the CKM and PMNS matrices
We use the standard parametrization of the CKM matrix
$V_{\rm CKM}=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-i\delta}\\ -s_{12}c_{23}-c_{12}s_{23}s_{13}e^{i\delta}&c_{12}c_{23}-s_{12}s_{23}s_{13}e^{i\delta}&s_{23}c_{13}\\ s_{12}s_{23}-c_{12}c_{23}s_{13}e^{i\delta}&-s_{23}c_{12}-s_{12}c_{23}s_{13}e^{i\delta}&c_{23}c_{13}\end{array}\right)\,.$ (C.1)
As four independent parameters we choose $V_{us}$, $|V_{cb}|$, $|V_{ub}|$ and the phase $\gamma$, so that
$s_{13}$ $=$ $|V_{ub}|\,\,,\quad s_{12}=\frac{V_{us}}{c_{13}}\,\,,\quad s_{23}=\frac{|V_{cb}|}{c_{13}}\,\,,\quad\delta=\gamma\,\,,$
$\quad c_{ij}$ $=$ $\sqrt{1-s_{ij}^{2}}\,\,.$ (C.2)
$V_{us}$ and $\gamma$ are set to $V_{us}=0.2253$ and $\gamma=64.6^{\circ}$. $|V_{cb}|$ and $|V_{ub}|$ are varied in the ranges spanned by the exclusive-inclusive measurements (69), with $|V_{cb}|_{exc}=(39.1\pm 0.5)\times 10^{-3}$, $|V_{cb}|_{inc}=(42.19\pm 0.78)\times 10^{-3}$, $|V_{ub}|_{exc}=(3.51\pm 0.12)\times 10^{-3}$, $|V_{ub}|_{inc}=(4.19\pm 0.17)\times 10^{-3}$ [5].
The standard parametrization is also used for the PMNS matrix
$U_{\rm PMNS}=\left(\begin{array}[]{ccc}k_{12}k_{13}&x_{12}k_{13}&x_{13}e^{-i\delta_{\rm PMNS}}\\ -x_{12}k_{23}-k_{12}x_{23}x_{13}e^{i\delta_{\rm PMNS}}&k_{12}k_{23}-x_{12}x_{23}x_{13}e^{i\delta_{\rm PMNS}}&x_{23}k_{13}\\ x_{12}x_{23}-k_{12}k_{23}x_{13}e^{i\delta_{\rm PMNS}}&-x_{23}k_{12}-x_{12}k_{23}x_{13}e^{i\delta_{\rm PMNS}}&k_{23}k_{13}\end{array}\right)\,,$ (C.3)
where $x_{ij}=\sin\theta_{ij}$ and $k_{ij}=\cos\theta_{ij}$. A global fit of the parameters in (C.3) gives values which depend on the assumed hierarchy among the neutrino masses [81]: for normal ordering
$\sin^{2}\theta_{12}$ $=$ $0.308^{+0.012}_{-0.011}\,\,,\quad\sin^{2}\theta_{23}=0.470^{+0.017}_{-0.013}\,\,,\quad\sin^{2}\theta_{13}=0.02215^{+0.00056}_{-0.00058}\,\,,$
$\delta_{\rm PMNS}$ $=$ $(212^{+26}_{-41})^{\circ}\,\,,$ (C.4)
for inverted ordering
$\sin^{2}\theta_{12}$ $=$ $0.308^{+0.012}_{-0.011}\,\,,\quad\sin^{2}\theta_{23}=0.550^{+0.012}_{-0.015}\,\,,\quad\sin^{2}\theta_{13}=0.02231^{+0.00056}_{-0.00056}\,\,,$
$\delta_{\rm PMNS}$ $=$ $(274^{+22}_{-25})^{\circ}\,\,.$ (C.5)
The global fit in [82] gives consistent results. In the numerical analysis we use the parameters for normal ordering (C.4).